OpenStudy (firejay5):

Factor each Polynomial. Show your work. Solve using Difference of Squares: A^2 - B^2 = (a -b) (a + b) Differences of Cubes: A^3 - B^3 = (A - b) (a^2 + ab + b^2) Sum of Cubes: A^3+ B^3 = (a + b) (a^2 - ab + b^2) 12. 5x^3 - y^3

OpenStudy (anonymous):

Well this is a difference of cubes\[5x ^{3}-y ^{3}\]=\[=(x(\sqrt[3]{5})-y)((x ^{2}5^{2/3})+(xy \sqrt[3]{5})+y ^{2})\]

OpenStudy (anonymous):

you just have to leave the 5 as a root 3

OpenStudy (firejay5):

just start easy on me

OpenStudy (anonymous):

your difference of cubes is misleading and not clear. it is actually

OpenStudy (firejay5):

I just started the problem duh and that's what I got so far

OpenStudy (anonymous):

\[(x ^{3}-y ^{3})=(x-y)(x ^{2}+xy+y ^{2})\]

OpenStudy (firejay5):

yes

OpenStudy (firejay5):

(5x - y) is what I have so far; what's next?

OpenStudy (anonymous):

the x^3 in this problem is 5x^3 when you cube root that, it becomes \[x(\sqrt[3]{5})\]

OpenStudy (anonymous):

you cube root the whole thing, not just the x

OpenStudy (firejay5):

I can't root the x in the calculator, I can cube the 5, but it gives me a decimal 1.709975947

OpenStudy (anonymous):

so you just leave it in cube root form like I did.

OpenStudy (firejay5):

we put it in the calculator like that, but we don't write it on our paper

OpenStudy (firejay5):

so where do I go from (5x - y)

OpenStudy (anonymous):

that's is the wrong step. you can't just ignore the 5 and only cube root the x and y

OpenStudy (firejay5):

what is it 5(x - y)

OpenStudy (anonymous):

why do you need to put that into the calculator?

OpenStudy (firejay5):

idk what ya mean

OpenStudy (anonymous):

you said you put it into the calculator and don't write it on paper right?

OpenStudy (firejay5):

cause we don't nee to write it on our paper; if it's in the calculator

OpenStudy (firejay5):

so what's my next step

OpenStudy (anonymous):

you need to cube root 5x^3 and y^3 put into the form (a-b)

OpenStudy (anonymous):

where 5x^ is a^3 and y^3 is b^3

OpenStudy (firejay5):

x and y ^3 will be x and y, so what about the 5